OpenStudy (anonymous):

y=4/7x+5/7 y=5/6x+5/7 solve the system of equations by using the substitution method

OpenStudy (anonymous):

for substitution you need to solve one of the equation for one of the variable, which they are already, so you can substitute 4/7x+5/7 for y in the second equation: 4/7x+5/7=5/6x+5/7 -5/7 -5/7 ------------------- 4/7x=5/6x and the only way that two different values multiplied by the same number can equal the same thing, is when x=0 so plug the x value into one of the equations: y=5/6x+5/7 and you get that y=5/7